# Optical-Cavity Manipulation Strategies of Conical Intersections Mediated Singlet Fission Systems

### v1

@@ Abstract
Figure 5: Real part of the SE (row (a)), GSB (row (b)), ESA (row (c)) contributions and the total 2DES spectrum (row (d)) of the two RD system for several population times $T_{w}$ indicated in the panels. The spectra are computed in the impulsive limit.
Brief summary. We have applied a numerically accurate mDA methodology to explore dynamic and spectroscopic responses of the CI-driven SF RD system strongly coupled to the cavity mode. Several microscopic SF mechanisms have been uncovered, viz. proximity of the CI-driven polaritonic wave packet to the Franck-Condon region, cavity-induced enchantment/weakening/suppression of SF, polaron/polariton decoupling, localization of the singlet-state population via engineering of the cavity-mode excitation, collective enhancement of SF.
Simulation and inspection of nonlinear femtosecond spectroscopic signals of the polaritonic CI-SF systems provides accurate information on the singlet-to-triplet transfer in real time. In particular, 2DES spectra of the RD system are dominated by the quartet of DPs and CPs that reveal the MP and UP bands, while the hybrid polariton-vibrational CPs contain information on the UP sideband. This clear separation of the polaritonic and vibrational peaks is a noticeable feature of the obtained spectra. In addition the simulated 2DES spectra exhibit multiple peak-cancellation effects, and their population-time evolution is accompanied by polariton-vibrational beatings.
We conclude that microcavities can open up new avenues for improving efficiency and stability of the CI-mediated SF, and ultrafast nonlinear spectroscopic signals – if interpreted with adequate theoretical support – can provide the real-time guidance on the SF reaction pathways. Yet, the polaritonic SF studies have not reached its maturity, and a lot of problems await efficient and practical solutions. The specifics are detailed below.
Open questions, challenges, and perspectives.
@@
System Hamiltonian. In excitonic systems, doubly-excited states are usually understood as the states in which two chromophores are simultaneously excited. This interpretation, however, is not always sufficient for SF aggregates, because polyatomic molecules comprising these aggregates possess dense manifolds of higher-lying excited electronic states and the minimal system supporting the SF reaction is not a single chromophore, but a dimer. Usually, higher-lying electronic states do not contribute to the SF per se (see the discussion below though). Yet, optical transitions between the lower-lying and higher-lying states affect dramatically ESA contributions to spectroscopic signals. Hence higher-lying states of individual molecules comprising SF dimers, aggregates, and materials have to be taken into account in realistic simulations of spectroscopic responses.
@@
It is useful to characterize all relevant electronic states and parameterize the SF Hamiltonians by using ab initio methods of quantum chemistry MT18 ; MT19 ; YiZhao21 ; Haibo21 ; Vendrell23 . In addition, ab initio methods can be applied to the entire system of molecules/aggregate+cavity. For example, quantum mechanics/molecular mechanics (QM/MM) simulations of Refs. Gerrit22 ; Gerrit22a treat up to 512 polyatomic chromophores in the cavity. Simulations based on the ab initio QED Rubio23 ; Koch21 , multiple spawning Martinez24 , and cavity Born-Oppenheimer Hartree-Fock Ansatz Kowalewski23 can also provide valuable insight into SF reactions in specific systems.
Quantum dynamic and spectroscopy. Up to now, accurate quantum simulations of up to two CI-carrying molecules/dimers in the cavity have been performed Gu1 ; sun1 ; sun2 . It is of great interest to push the simulations towards higher number of molecules/dimers $N_{M}$ and explore how CIs affect cooperativity, e.g., to check the conventional $\sim\sqrt{N_{M}}$ scaling of the effective system-cavity coupling Huo22 . For this to achieve, few-molecule models with symmetry-optimized Hamiltonians JYZ23a as well as multi-configuration time-dependent Hartree (MCTDH) MT18 ; MT19 ; Vendrell19 ; Vendrell20 , tensor-network Vidal18 ; Keeling22 , and mDA methods look promising. Recently emerging emulations of CIs on quantum simulators CIsimulations23a ; CIsimulations23b ; Schoelkopf23 may also be extended in the future to cavity-CI systems.
Disorder is of paramount importance for realistic modeling of CI-driven cavity-mediated SF systems, taking into account that defects/disorder can induce CIs in semiconducting molecular materials Chuntonov22 and affect significantly polaritonic transport Schachenmayer22 ; HTC . Temperature is essential, e.g., for activating CI-mediated SI in rubrene Miyata and for catalysing reactions in polaritonic systems Campos . The mDA protocol based on the thermo field dynamics (TFD) BG21 was recently applied to explore dynamics of the pristine RD skw ; Sun20 and the Holstein-Tavis-Cummings model Hou23 at finite temperature.
Apart from mDA, MCTDH and variants of tensor-network methods are especially favorable for performing fully-quantum simulations of realistic polaritonic CI-driven SF systems, because they permit the efficient evaluation of nonlinear spectroscopic signals Santoro23 ; GB_JCTC and – by using the TFD methodology BG21 and the recipe of Ref. GB21 – finite temperature and static disorder can also be taken care of in the (relatively) computationally inexpensive manner.
@@
An extra challenge is to identify spectroscopic signatures of polaritonic effects in nonlinear spectroscopic signals of polaritonic CI-driven SF systems. Hence accurate theoretical simulations and interpretations of the existing Takahashi ; Watanabe20 ; Chergui20 and forthcoming experiments are demanding but rewarding.
@@
Microscopic mechanisms behind polaritonic CI-driven SF. There exist several methods that can potentially enhance the SF yield. The conventional practices are based on the adjustment of the cavity mode for achieving resonance and strong coupling of the LP or UP states with the triplet state or on the construction of the cavity-assisted SF pathways. On the other hand, the geometric-phase effects may also be harnessed for engineering SF in CI-driven polaritonic systems. The controlling mechanisms may be implemented through collective topological polaritonic phases JYZ23 that are akin to those in topological metamaterials Alu23 . It is also worthwhile to explore interfacing of light-induced CIs in cavities Vibok22a ; Vibok22b ; Huo21 with intrinsic CI systems, as their combination may facilitate fine-tuning of the multiple CI landscapes VendrellVibok .
@@
Apart from the aforementioned cavity-related mechanisms, there exist other handles for tweaking SF, that are based on the optimization of the bias voltage Xiong18 , temperature Miyata ; Xiong15 ; Wasielewski21 , carrier frequency BY22 and intensity Asahi16 ; Bardeen18a of the excitation pulse, or even static disorder Chavez ; Disorder24 .
@@
Yet, it is now always clear whether the coupling to the cavity (perhaps in combination with other methods mentioned above) enhances or suppresses the SF yield. It is promising, therefore, to interface the existing protocols for the dynamic simulations of polaritonic CI-driven SF systems with quantum control routines Control1 ; Control2 , in order to find a robust way to comprehensive optimization of SF in cavities.
I TOC Graphic

### v2

@@ III.3 Spectroscopic signatures of Polaritonic CIs
Figure 8: Real part of the SE (row (a)), GSB (row (b)), ESA (row (c)) contributions and the total 2DES spectrum (row (d)) of the two RD system for several population times $T_{w}$ indicated in the panels. The spectra are computed in the impulsive limit. Adapted from sun2 . Copyright American Chemical Society.
Here we consider electronic two-dimensional (2D) spectra, because electronic 2D spectroscopy (2DES) is one of the most powerful and information-rich third-order spectroscopic techniques (see Scholes22 and references therein). Qualitatively, 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ can be envisaged as a $T_{w}$-dependent 2D $(\omega_{\tau},\omega_{t})$ maps, which reveal correlations between the states of the system excited at a frequency $\omega_{\tau}$ at time zero and probed at a frequency $\omega_{t}$ at time $T_{w}$. 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ simulated for the two RD system strongly coupled to the cavity mode are presented in Fig. 8 for different population times $T_{w}$. From top to bottom, shown are the SE (row (a)), GSB (row (b)), and ESA (row (c)) contributions, as well as the total signal (row (d)). The reader is referred to Appendix B for more derivation details on the various nonlinear response functions within the framework of the multi-D2 Ansatz.
The SE contribution (row (a)) possesses two pronounced diagonal peaks (DPs) and two cross-peaks (CPs) (similar situation takes place in the purely polaritonic dimer system Bondarev13 ). The DP1 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and DP2 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) reveal the middle polariton (MP) and the UP bands, respectively. The low-intensity LP band cannot be reliably identified in the SE 2DES spectrum, because 2DES peak intensities are proportional to the product of four relevant transition dipole moments, in contrast with absorption spectra, which are proportional to the product of two transition dipole moments (see the discussion in Ref. Jindra15 ). The CP1 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and CP2 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) quantify the coupling and polaritonic coherence between the states MP and UP. A weak CP ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) at $T_{w}=16$ fs reveals the phonon sideband. The shape and intensity of the DPs and CPs depend significantly on the waiting time $T_{w}$. The DPs at $T_{w}=0$ are predominately elongated perpendicular to the main diagonal, but their elongation turns into parallel at longer population times. A certain similarity of the DPs at $T_{w}=0$ and 32 fs shows a significant contribution of the polaritonic beating with a period of $\sim 32\leavevmode\nobreak\ {\rm fs}$ revealing the energy difference of $0.132\,{\rm eV}$ between the states MP and UP. The two negative CPs are elongated parallel to the main diagonal. Similar to the DPs, the CPs are also significantly modulated by the MP-UP energy-gap beating. The overall decreases of the intensity of the SE contribution with population time mirrors the CI-mediated population transfer from the bright ${\rm S}_{1}^{(j)}$ states to the dark ${\rm TT}^{(j)}_{1}$ states and reflects thereby the SF dynamics.
The GSB contribution is shown in panels (b) of Fig. 8. It reveals the same pairs of the polaritonic DPs and CPs and, additionally, a pair of weaker polariton-phonon CPs, ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$). In comparison with the SE peaks, their GSB counterparts look rather static. Nevertheless, the GSB contribution reveals the vibrational wave-packet motion in the electronic ground state, which is clearly manifested in the peak evolutions (see Ref. sun2 ).
The ESA contributions are depicted in panels (c) of Fig. 8. The SE and ESA contributions to 2DES spectra reveal photophysical processes in the manifold $N=1$ skw20a . However, the two contributions employ different reporter states for delivering the same information: the SE employs the electronic ground state $N=0$, while the ESA employs the doubly-excited manifold $N=2$. Hence shapes and evolutions of predominantly positive SE contributions and predominantly negative ESA contributions are quite similar. Nevertheless, positions of the DPs and CPs in the SE and ESA spectra are slightly shifted with respect to each other. At $T_{w}=0\leavevmode\nobreak\ {\rm fs}$, for example, the (nearly) diagonal ESA peaks are located at ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}\approx 2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.416\,{\rm eV}$). All ESA peaks are elongated parallel to the $\omega_{t}$ axis, due to a high density of states in the manifold $N=2$. The peak ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$), which revels transitions from the MP state to the states of the doubly-excited manifold, has the highest intensity. Note also that the relative separation between the DPs and CPs along the $\omega_{t}$ axis changes with $T_{w}$: cf. the spectra at 0 fs with those at 16 fs. This modulation is caused by the superposition of polaritonic and purely vibrational tuning-mode oscillations. This reflects a higher complexity of the ESA contribution, which involves polaritonic states with $N=1$ and 2.
Finally, the total 2DES signal $S(\omega_{\tau},T_{w},\omega_{t})$ is shown in Fig. 8(d). Since positions of the DPs and CPs in the SE, GSB and ESA contributions nearly coincide, the total 2DES spectrum exhibits positive, negative and dispersed peak shapes, the amplitudes of which are determined by subtle cancellation effects in the superposition of positive SE and GSB contributions and negative ESA contribution. The overall structure of the total 2DES spectra is dominated by the quartet of polaritonic DPs and CPs. In addition, a number of polariton-vibrational peaks around $\omega_{\tau}\,\mathrm{or}\,\omega_{t}=$ 2.6 eV is seen at different $T_{w}.$ It is remarkable that these hybrid polariton-vibrational peaks are more pronounced in the total signal than in the SE, GSB, and ESA contributions – owing to the aforementioned cancellation effects. The DPs are positive while the CPs exhibit unusual asymmetry: the upper CP is always positive, but the spectral structures around the lower CP can be negative (see the signals for $T_{w}=0$ and 16 fs).
IV Simulation of polaritonic systems: unsolved problems, challenges, and outlooks
IV.1 Model and Hamiltonian
While the essential features of polariton dynamics can be effectively captured using the HTC Hamiltonian, the study of more practical SF-related photophysical processes has multifaceted significance and application value. The direct ${\rm S}^{(1)}\rightarrow{\rm TT}^{(1)}$ CI-driven SF is one of the possible mechanisms. In specific systems, charger transfer, locally-excited, doubly-excited, and multi-excitonic states facilitate the SF process, determine its pathways, and govern its efficiency MT18 ; MT19 . Furthermore, the coupling between these states can be of different nature, namely a dipolar coupling (which corresponds to the avoided crossing between the corresponding electronic/excitonic PESs) or a vibrational-mode mediated coupling (which corresponds to the CI between the corresponding electronic/excitonic PESs). Obviously, embedding a multi-mode, multi-state system into a microcavity produces a cascade of coupled polaritonic states and PESs governing the SF process. If several SF systems are embedded in one cavity, the situation becomes even more complex. It is essential though that all these polaritonic states directly or indirectly contribute to the coupling between the parent state ${\rm S}^{1}$ and the product state ${\rm TT}^{1}$. Hence, their net effect can be treated by introducing an effective ${\rm S}^{(1)}$-${\rm TT}^{(1)}$ coupling strength. We thus anticipate that the dynamical SF effects described here remain qualitatively valid if a large number of relevant electronic states is included.
A proper characterization of cavity-mediated SF requires a careful selection of the relevant excitonic states and mechanisms of their couplings, notably for a large number $N_{M}$ of SF species embedded in the cavity. In this context, it is appropriate to mention a recent discussion on the possibility to reach a strong-coupling regime in the cavities hosting a large number of species. Liu et al. pinpointed significant challenges in modifying photophysical dynamics in an optical microcavity, necessitating a molecular system with long excited-state lifetimes or architectures that enhance exciton-photon coupling strengths. These requirements are stipulated by the existence of the (quasi)degenerate manifold of dark “reservoir” states (e.g., triplets) as well as the slow conversion from reservoir states to cavity polaritons liu1 . In other words, the results of Liu et al. liu1 as well as those of Takahashi and Watanabe Watanabe20 seem to demonstrate that a good portion of the singlet-state population bypasses the correlated triplet state and goes to the lower dark polaritonic states in the strong-coupling regime. On the other hand, a recent report by the Chergui group Chergui20 does not support this scenario. There appears to be a consensus that this challenge can be addressed through a careful and strategic redesign of the entire cavity QED system. Menon24 .
@@ IV.1 Model and Hamiltonian
In excitonic systems, doubly-excited states are usually understood as the states in which two chromophores are simultaneously excited. This interpretation, however, is not always sufficient for SF aggregates, because polyatomic molecules comprising these aggregates possess dense manifolds of higher-lying excited electronic states and the minimal system supporting the SF reaction is not a single chromophore, but a dimer. Usually, higher-lying electronic states do not contribute to the SF per se (see the discussion below though). Yet, optical transitions between the lower-lying and higher-lying states affect dramatically ESA contributions to spectroscopic signals. Hence higher-lying states of individual molecules comprising SF dimers, aggregates, and materials have to be taken into account in any realistic simulation of spectroscopic responses.
Taking into account the multimode (essentially, quasicontinuum) spectral densities of the microcavities Demler20 ; Gerrit22a ; Ribeiro22 and finite cavity-mode lifetimes Huo24 may be essential for a proper description of the polaritonic dynamics in SF systems. Thus, a single continuous-mode description of the system-cavity coupling may be insufficient or even inadequate. However, the vertical excitation energy of a single dimer is around 2.2 eV. In this case, an effective single-mode description can be adopted for multi-mode cavities as only those modes in resonance with this transition would substantially contribute to the polaritonic SF. Similar effects have been explicitly demonstrated, e.g., in Ref. GB24 where the fine (400 modes) description of the photonic spectral density was adopted.
It is crucial that all the additional states and specific inter-state couplings mentioned above can be effectively handled using the multi-D2 Ansatz when required. Moreover, the methodology employing multi-D2 Ansatz is efficient for arbitrary time-dependent Hamiltonians, non-Hermitian Hamiltonians, and models of multispecies, multimode bosonic baths zhao2 ; zhao1 . Hence, the multi-D2 Ansatz is a method of choice for handling quasicontinuous spectral densities of the cavity modes, finite lifetimes of the cavity/plasmonic modes, and fine discretization schemes of the cavity modes.
@@
Characterizing all relevant electronic states and parameterizing the SF Hamiltonians using ab initio quantum chemistry methods proves to be highly beneficial. MT18 ; MT19 ; YiZhao21 ; Haibo21 ; Vendrell23 . In addition, ab initio approaches can also be applied to the entire systems of molecules/aggregates including the cavity. For example, quantum mechanics/molecular mechanics (QM/MM) simulations of Refs. Huo ; Gerrit22 ; Gerrit22a treat up to 512 polyatomic chromophores in a cavity. Simulations such as those based on the ab initio QED Rubio23 ; Koch21 , the multiple spawning method Martinez24 , and the cavity Born-Oppenheimer Hartree-Fock Ansatz Kowalewski23 can also provide valuable insights into dynamics of SF reactions in specific systems.
IV.2 Dynamics and spectroscopy
Usually, accurate quantum simulations of up to two CI-carrying molecules/dimers in a cavity are performed Gu1 ; sun1 ; sun2 . Vendrell et al., however, have simulated up to seven pyrazine molecules in a cavity, with each molecule having four modes and three states. They conducted a detailed analysis of the coupling between the cavity and the resonances of $S_{0}/S_{2}$ as well as $S_{1}/S_{2}$ Vendrell . It is of great interest to push the simulations towards higher number of molecules/dimers $N_{M}$ and explore how CIs affect cooperativity, e.g., to check the conventional $\sim\sqrt{N_{M}}$ scaling of the effective system-cavity coupling TAVIS ; Huo22 .
To achieve this aim, a combination of advanced theoretical models and computational techniques appears promising. Few-molecule models with symmetry-optimized Hamiltonians JYZ23a provide a simplified yet effective framework for understanding key interactions. Complementing these models, the multi-D2 Ansatz, as utilized here zhao1 ; zhao2 , offers a powerful variational approach to accurately capture the dynamics of multi-species, multi-modes systems with a large number of DOFs. In particular, the HTC model containing dozens of qubits is readily tackled with the multi-D2 Ansatz technique HTC . Additionally, the multi-configuration time-dependent Hartree (MCTDH) method MT18 ; MT19 ; Vendrell19 ; Vendrell stands out as a robust tool for exploring high-dimensional quantum dynamics, particularly in systems where vibrational and electronic degrees of freedom are strongly coupled. The multi-layer extension of MCTDH has been carefully compared with the multi-$\rm D_{2}$ Ansatz method by Saalfrank and coworkers in a very recent study of vibrational relaxation dynamics at surfaces Fischer . While both rely on the time-dependent variational principle to arrive at an optimal wave function via solving highly nonlinear coupled equations of motion for variational parameters, the mDA method was found to be much more efficient (15,000s on a modern laptop) than the multilayer MCTDH (54,000s on a modern workstation). Furthermore, another potentially useful approach is the so-called time evolving density operator with orthogonal polynomials algorithm (TEDOPA) Prior2010 ; Chin2011 . Through a basis change of the Hamiltonian, the TEDOPA method transforms the environment into a chain of harmonic oscillators with only local nearest-neighbor interactionsGuimares ; Tamascelli . By utilizing Matrix Product States (MPS) for simulating the dynamics of one-dimensional quantum systems, TEDOPA effectively circumvents the otherwise unmanageable scaling associated with system size. Yet, the multi-$\rm D_{2}$ Ansatz method seems to be more flexible; for example, it has been applied very recently to the two-dimensional t-J model shen2024 . Furthermore, the multi-$\rm D_{2}$ Ansatz method can also address systems with long-range or strong interactions, as well as those involving entanglement. These challenges typically require a higher bond dimension in the MPS, leading to a long simulation time with TEDOPA. It is worth mentioning that the tensor-network techniques Vidal18 ; Keeling22 have emerged as an effective approach for handling the computational challenges associated with large-scale quantum systems.
Recently emerging emulations of CIs on quantum simulators CIsimulations23a ; CIsimulations23b ; Schoelkopf23 may also be extended in the future to cavity-CI systems. Unfortunately, the development of QM/MM approaches and numerically accurate fully quantum methods has largely progressed in isolation, with limited integration or cross-validation. This separation creates a significant gap in leveraging the strengths of both methodologies. Fully quantum methods provide high-precision insights into the dynamics of SF systems but are computationally intensive. In contrast, QM/MM methods offer a more computationally feasible framework by combining quantum-level detail with classical mechanics approximations, making them suitable for simulating larger systems or processes. Given these complementary strengths, it is highly desirable to bridge this gap by systematically validating QM/MM methods against fully quantum calculations. Such a validation process would not only enhance the reliability of QM/MM approaches but also enable their broader application in scenarios where fully quantum simulations are computationally prohibitive.
Disorder is of paramount importance to any realistic modeling of CI-driven cavity-mediated SF systems, as defects and disorder in semiconducting molecular materials can give rise to the formation of CIs Chuntonov22 and affect polaritonic transport substantially Schachenmayer22 ; HTC . Temperature adjustment is essential, e.g., for activating CI-mediated SI in rubrene Miyata and for catalysing reactions in polaritonic systems Campos . The multi-D2 Ansatz protocol based on the TFD BG21 was recently applied to explore dynamics of the pristine RD skw ; Sun20 . Subsequently, Hou et al. utilized this approach to obtain the finite-temperature dynamics and the absorption spectra of HTC model HouE24 , as referred to in Sect. II.3. Looking forward, combining the multi-D2 Ansatz with the TFD method, exploration of cavity-controlled SF processes at finite temperatures is anticipated to uncover new mechanisms, including the impact of vibrational polarization on the SF process induced by strong vibrational coupling, the influence of nonlinear coupling terms (such as the Kerr effect) within the cavity, and the role of multi-mode cavities. These factors are expected to provide deeper insights into the interplay between cavity dynamics and SF processes, potentially revealing novel pathways for enhancing efficiency and control. Apart from multi-D2 Ansatz, MCTDH and variants of tensor-network methods are also well-suited for conducting fully quantum simulations of realistic polaritonic CI-driven SF systems, as they enable efficient evaluation of nonlinear spectroscopic signals Santoro23 ; GB_JCTC and – by using the TFD method BG21 and the recipe of Ref. GB21 – effects of finite temperatures and static disorder can also be taken care of in a (relatively) computationally inexpensive manner.
@@ IV.2 Dynamics and spectroscopy
Yet a microscopic and fully quantum description of spectroscopic experiments has to await future developments: The bottleneck is accounting for the large number of system and cavity DOFs and necessity to perform simulations on relatively long (picosecond to microsecond) timescales. An extra challenge is to identify spectroscopic signatures of polaritonic effects in nonlinear spectroscopic signals of polaritonic CI-driven SF systems. Hence, accurate theoretical simulations and interpretations of the existing Takahashi ; Watanabe20 ; Chergui20 and forthcoming experiments are demanding but rewarding.
IV.3 Mechanisms of polaritonic CI-driven SF
There exist several methods that can potentially enhance the SF yield. Conventional practices are based on the adjustment of the cavity mode for achieving resonance and strong coupling of the LP state with the triplet state. Moreover, Vendrell et al. discovered the possibility of using the UP state as an entry door for the cavity-assisted SF pathways in a rigid system, which leads to a much higher yield Vendrell3 .
@@ IV.3 Mechanisms of polaritonic CI-driven SF
On the other hand, the geometric-phase effects may also be harnessed for engineering SF in CI-driven polaritonic systems. The controlling mechanisms may be implemented through collective topological polaritonic phases JYZ23 that are akin to those in topological metamaterials Alu23 . It is also worthwhile to explore interfacing of light-induced CIs in cavities Vibok22a ; Vibok22b ; Huo21 with intrinsic CI systems, as their combination may facilitate fine-tuning of the multiple CI landscapes VendrellVibok .
@@
Apart from the aforementioned cavity-related mechanisms, there exist other handles for tweaking SF, which are based on the optimization of the bias voltage Xiong18 , temperature Miyata ; Xiong15 ; Wasielewski21 , carrier frequency BY22 and intensity Asahi16 ; Bardeen18a of the excitation pulse, or even static disorder Chavez ; Disorder24 .
@@
Yet, it is not always clear whether the coupling to the cavity (perhaps in combination with other methods mentioned above) enhances or suppresses the SF yield. It is promising, therefore, to interface the existing protocols for dynamic simulations of polaritonic CI-driven SF systems with quantum control routines Control1 ; Control2 , in order to find a robust pathway to comprehensive optimization of SF in cavities.
V Conclusions
Integrating molecular materials into optical cavities has ushered in a new era of controlling quantum dynamics in complex systems. By coupling molecular excitations with quantized light fields, cavity environments reshape exciton energies, open new reaction pathways, and offer routes to surpass traditional photoconversion limits. This Perspective highlights recent developments in fully quantum simulations to understand CI-driven SF process in an optical cavity. A numerically accurate Davydov trial state, the multi-D2 Ansatz, is employed in a time-dependent variational framework to explore dynamic and spectroscopic responses of the TC model, the HTC model, and the CI-driven SF RD system strongly coupled to a cavity mode. For the disordered TC model, time evolution of the photon population and the optical absorption spectra are also calculated by using the Green’s function method to corroborate numerically accurate results of the multi-D2 Ansatz. The impacts of temperature, static disorder, photon loss, and qubit-phonon coupling on the dynamics and absorption spectra of the HTC model are investigated by combining the multi-D2 Ansatz combined with the TFD method.
Several microscopic SF mechanisms have been uncovered, viz. proximity of the CI-driven polaritonic wave packet to the Franck-Condon region, cavity-induced enchantment/weakening/suppression of SF, polaron/polariton decoupling, localization of the singlet-state population via engineering of the cavity-mode excitation, collective enhancement of SF.
Simulation and inspection of nonlinear femtosecond spectroscopic signals of the polaritonic CI-SF systems provides accurate information on real-time singlet-to-triplet transfer. In particular, 2DES spectra of the RD system are dominated by the quartet of DPs and CPs that reveal the MP and UP bands, while the hybrid polariton-vibrational CPs contain information on the UP sideband. This clear separation of the polaritonic and vibrational peaks is a noticeable feature of the spectra obtained. In addition, simulated 2DES spectra exhibit multiple peak-cancellation effects, and their population-time evolution is accompanied by polariton-vibrational beatings.
The findings emphasize the transformative role of polaritonic states in controlling SF dynamics, highlighting key factors such as cavity mode tuning, Rabi splitting, and electron-phonon interactions. Moreover, the study underscores the importance of integrating advanced simulation techniques, such as the multi-D2 Ansatz, to capture the complex interplay of electronic, vibrational, and photonic states in many-body systems.
Strictly speaking, the level of description of the CI-driven cavity-enhanced SF systems adopted in this perspective and the ensuing effects outlined above pertain to so-called picocavities, in which the strong-coupling regime can be achieved for a few embedded molecules Bhuyan . To validate these predictions, it will be worthwhile to perform ultrafast spectroscopic (perhaps polarization-sensitive) experiments on a series of well characterized SF systems (such as dimers, trimers, and tetramers) in solution and in picocavities. Furthermore, the effects akin to those demonstrated in this Perspective can likely be found in more realistic CI-driven SF systems containing a large number of SF species strongly coupled to multi-mode cavities.
It is concluded that microcavities can open up new avenues for improving efficiency and stability of the CI-mediated SF, and ultrafast nonlinear spectroscopic signals – if interpreted with adequate theoretical support – can provide a real-time guidance on the SF reaction pathways. Yet, polaritonic SF studies have not reached the level of maturity, and many problems await efficient and practical solutions. Looking ahead, the combination of fully quantum simulations, insightful modeling tools, and experimental innovations can guide us toward next-generation SF materials and cavity designs. By addressing these unsolved problems and challenges, researchers have the potential to leverage CIs and cavity architectures for surpassing conventional efficiency limits in solar energy harvesting and paving the way for groundbreaking innovations in next-generation photovoltaics and other optoelectronic applications.